OpenStudy (anonymous):

How do type out the cube root of 3

OpenStudy (turingtest):

in LaTeX you mean?

OpenStudy (anonymous):

you could just put parantheses around it and raise it to the (1/3) power

OpenStudy (jiteshmeghwal9):

``` sqrt[3]{} ```

Parth (parthkohli):

sqrt[3]{3}

Parth (parthkohli):

threeception. \[\sqrt[3]{\sqrt[3]{\sqrt[3]{3}}}\]